### v4

# Dynamically Augmented CVaR for MDPs

Abstract
This paper studies optimization of Conditional Value-at-Risk (CVaR) for Markov Decision Processes (MDPs) with finite state and action sets. It introduces the Dynamically augmented CVaR (DCVaR) risk objective function and provides an algorithm for its optimization. This paper investigates a specially defined Robust MDP (RMDP), in which the state space is augmented with the tail risk level. This RMDP, which we call the Dynamically augmented RMDP (DRMDP), was introduced to the literature for calculations of optimal CVaR values by value iteration more than ten years ago, but, as was understood later, these value iterations compute lower bounds of minimal static CVaRs. DCVaR is defined as a time consistent version of the static CVaR, and it is a lower bound of the static CVaR. It also can be considered as a dynamic version of the nested CVaR. This paper provides an algorithm constructing a policy optimizing DCVaR of total discounted costs. The correctness of this algorithm is proved by studying a special mass transfer problem. The results on RMDPs needed for this paper are provided in the appendix.
Keywords: Markov decision process, conditional value at risk, optimal policy, algorithm
1 Introduction
Conditional Value-at-Risk (CVaR), also sometimes called Average Value-at-Risk (AVaR), is one of the most important and popular risk measures. For dynamic problems, it is possible to consider the so-called static CVaR, when the CVaR value is defined for each policy, and the goal is to find a policy optimizing this value. However, there are problems with this approach dealing with computational complexity of finding optimal policies and time inconsistency of static CVaRs. The issue of time consistency of risk measures is discussed in several works, including [22, 34, 35, 37, 38, 39]. An alternative approach is to consider the nested CVaR, which is defined via a Robust Markov Decision Pprocess (RMDP) whose state space coincides with the original state space of the problem. Another approach, which is also often used in practice, is to consider an RMDP, whose states are pairs consisting of the original state augmented with the risk level. In other words, the state is dynamically augmented by a number between 0 and 1 representing the tail risk level.
The theory for static CVaR was developed by B¨auerle and Ott [2], and nested CVaR was studied by Ruszczy´nski [33] and Shapiro [39]. Using the CVaR decomposition theorem by Pflug and Pichler [29], Chow et al. [4] introduced an RMDP with the risk-augmented states for computing optimal CVaR values by value iteration. Approximately 10 years later Hau et al. [16] proved that the result of the value iteration may not be equal to the optimal value of the static CVaR, and it is a lower bound for the minimal static CVaR.
This paper, which deals with problems with finite state and action sets, introduces a Dynamically augmented Conditional Value-at-Risk (DCVaR) defined by using the RMDP introduced in Chow et al. [4], which we call the Dynamically augmented RMDP (DRMDP). The main result is Algorithm DCVaR for constructing an optimal policy minimizing the DCVaR value for a given initial tail risk level.
The first group of the results in this paper establishes the relation between the static CVaR values and the DRMDP. First we observe that there is an optimal nonrandomized policy minimizing static CVaR, and, for a nonrandomized policy for a Markov Decision Process (MDP), its static CVaR is equal to the worst possible expected outcome if the Decision Maker (DM) plays this policy in the DRMDP. This observation clarifies the gap described in Hau et al. [16] and illustrates time inconsistency of the static CVaR for MDPs because, in order to implement this worst possible outcome, the second player in the DRMDP, whom we call Nature, may have to play a policy, which knows future decisions of the DM. A more natural assumption is that Nature plays its optimal policy, and this assumption leads to the definition of the DCVaR provided in this paper. This definition implies that value iterations proposed in Chow et al. [4] converge to minimal values of DCVaR.
The second group of the results in this paper deals with formulating Algorithm DCVaR for constructing optimal DCVaR policies. The proof that this algorithm constructs optimal DCVaR policies is based on the structure of solutions to a special mass transfer problem describing Nature’s optimal decisions.
The optimality equations for the DCVaR is relevant to the optimality equation for the nested CVaR, and the nested CVaR can be interpreted as the DCVaR with the fixed tail risk level, while optimal DCVaR policies make decisions for variable tail risk levels. The tail risk levels for optimal DCVaR policies depend on the initial tail risk level and gains and losses incurred up to the current time epoch. The DM knows only the initial risk level and does not observe risk levels at later stages, but Algorithm DCVaR implicitly estimates current risk levels.
Classic methods for MDPs deal with risk-neutral objectives. In addition, there is a significant literature on risk-sensitive optimization initiated by Howard and Matheson [19] dealing mostly with optimization of expectations of utility functions rather than with optimization of expected total costs or average costs per unit time. Pioneering work by Markowitz [26] on mean-variance optimization for portfolio selection influenced research on MDPs with mean-variance criteria [11, 42], but variance of the total costs is not a convenient characteristic to deal with for sequential problems, and it is not a good risk measure, because variance is affected by losses and gains in the same way. CVaR is an attractive alternative to variance both because of its practical importance and rich mathematical properties; see [30, 31, 32, 40]. CVaR is a risk measure that has gained popularity in various applications including finance. It has the ability to safeguard a DM from the tail events by focusing on an average of the largest losses and by doing so it provides a more comprehensive view for risk management than threshold-based risk measures such as Value-at-Risk(VaR). CVaR is also known as expected shortfall in financial literature and has great importance in terms of financial regulation.
We summarize the organization of this paper and its main results. Preliminary information on CVaR and MDPs is provided in Section 2. In Section 3 we prove the existence of a nonrandomized policy minimizing CVaR of the total discounted costs. Then we define the DRMDP as an RMDP in which the DM chooses actions, and Nature assigns tail risk levels. States of the DRMDP are states of the original MDP augmented with tail risk levels. Policies for the original MDP are particular policies of the DM in the DRMDP when the DM does not observe past and current tail risk levels, except for the initial one, and we call these policies risk-independent. We show that for a nonrandomized risk-independent policy played by the DM, CVaR of total discounted costs is equal to the worst expected total discounted costs the DM can incur by playing this policy in the DRMDP. This result illustrates time inconsistency of the static CVaR and implies that the minimal static CVaR value for the MDP is equal to the minimal value for the class of nonrandomized risk-independent policies in the DRMDP. This equality illustrates a positive gap between the optimal static CVaR value and the limit of value iterations. This gap was discovered in Hau et al. [16]. Section 4 defines a DCVaR, describes its properties, and introduces a DRMDP1, which is obtained from the DRMDP by modifying one-step costs and transition probabilities. The payoff and value functions for the DRMDP1 are equal to these functions for the DRMDP multiplied by tail risk levels. The sets of optimal policies for the DRMDP and DRMP1 coincide if the initial tail risk level is positive. The advantage of the DRMDP1 is that its value function is concave in the tail risk level.
Section 5 introduces the algorithm for constructing a nonrandomized optimal policy minimizing DCVaR, which was announced in [5]. Sections 6–8 deal with proving that this algorithm constructs an optimal policy. Section 6 describes the properties of the mass transfer problem relevant to optimal solutions for Nature. Section 7 describes the properties of optimal values and properties of the sets of optimal actions. Section 8 proves correctness of the algorithm. Section 9 provides extensions to random cost. Appendix A provides results on RMDPs used in this paper.
2 Preliminaries: CVaR and MDPs
Conditional Value at Risk (CVaR) is a risk measure widely used in engineering and finance. For a random variable Z defined on a probability space its CVaR for a tail risk level is a conditional expectation of the tail. One of several equivalent formal definitions of CVaR is
and the Value at Risk is
where is the distribution function of In addition, if this expectation exists, and where by definition; see [30, 31, 32] and [40, Chapter 6] for additional details regarding CVaR.
It is well known that CVaR can be equivalently written in other forms. For
where for a number CVaR can also be represented using its dual representation
where , and is the set of random variables on the probability space defined as and this set is called the dual CVaR risk envelope. In this paper, we sometimes write a constant instead of a function equal to this constant, and comparisons between random variables are usually understood -a.s.
Let be a sub -algebra of that is, is a -algebra on and A -measurable random variable is called a random tail risk level. Then the conditional CVaR given the -algebra and a random tail risk level is defined [29, Definition 17] as the -measurable random variable
where the essential supremum is taken with respect to the conditions probability and is the set of nonnegative -measurable random variables on such that and According to the CVaR decomposition theorem by Pflug and Pichler ([29], Lemma 22),
This paper deals with minimizing CVaR of a random variable defined as
where are random variables, and either and or and The constant is called a discount factor. In particular, this paper deals with the situation when the stochastic sequence can be controlled. This means that the probability measure can depend on a policy chosen by a DM, and the goal is to choose the best policy. The theory of Markov Decision Processes (MDPs) provides a modeling framework for such problems.
An MDP is a model of a natural family of controlled stochastic sequences. It is a tuple where is a set of states, is a set of actions, and for each state there is a nonempty set of available actions is a real-valued cost function, and is a transition probability. The time is discrete, and, if at some time instance an action is selected at a state , then the system moves to the next state according to and the cost is collected, allowing the cost to depend on the next state. In this paper we always assume that and are finite sets.
Let where be the set of histories up to epoch and is the set of trajectories. This set is endowed with the -algebra which is the countable product of -algebras consisting of all subsets of the finite sets
A policy is a sequence of transition probabilities from the sets of finite histories to the sets of actions such that for all Let be the set of all policies. According to the Ionescu Tulcea theorem, an initial state and a policy define a probability on the space of trajectories where are transition probabilities from to ana is a transition probability from to which can be viewed as transition probabilities from to Expectations with respect to probabilities are denoted by If each probability is concentrated at one point, the policy is called nonrandomized. A nonrandomized policy is defined by a sequence of mappings such that for all and for all A nonrandomized policy is called Markov if each time the choice of a decision depends only on the current time and state. A Markov policy is defined by a sequence of mappings such that for all and for all A Markov policy is called deterministic if decisions do not depend on the time parameter. A deterministic policy is defined by a function such that for all Let be a constant discount factor.
An objective criterion is a real-valued function on In general, it can take infinite values, but in this paper we consider only real-valued objective criteria. The function where is called the value function. A policy is optimal for the initial stated if A policy is called optimal if it is optimal for all initial states In the literature these definitions are usually applied to risk-neutral objectives, but they can be applied for arbitrary objective functions [7] with values in
For a finite horizon the total discounted cost mentioned in (2.1) is the random variable
where is the terminal cost, and for the infinite-horizon
where if and if . In fact, for the results of this paper holds for nonnegative real-valued discount factors
The risk-neutral approach is to consider the expected total discounted cost For a tail risk level and a finite horizon or an infinite horizon let us consider the objective function where , and is the random variable defined on the probability space in formulae (2.2) or (2.3). The value is sometimes called a static CVaR. Computing optimal policies minimizing the static CVaR is a hard problem [2, Remark 5.1].
Chow et al. [4] changed the original problem formulation with the state space to the formulation when the states of the problem consist of state-risk pairs where is the original state and is the tail risk level. In other words, the state space is augmented in [4] with the tail risk level, and the new state space is Motivated by the Pflug-Pichler CVaR decomposition theorem [29], Chow et al. [4] introduced the Robust MDP (RMDP) with the state space which is formally defined in Section 3 and called the Dynamically augmented RMDP (DRMDP). Infinite horizon was considered in [4], and a finite horizon version was considered in [23], where it was assumed that the DM knows tail risk levels at all steps. Hau et al. [16] showed that the value of DRMDP is a lower bound of the minimal CVaR values for the original MDP. In Section 4 we introduce the Dynamically augmented CVaR (DCVaR) objective function, which is a lower bound of CVaR, and the value of DRMDP is equal to the minimal value of DCVaR.
In the conclusion of this section, we recall two concepts for MDPs relevant to more general objective criteria than expected total costs: MDPs with general objective functions and abstract dynamic programming. According to [7], a general objective function is a function such that, if for and then This definition is consistent with the notion of a law-invariant risk measure [29, 40].
3 Static CVaR and Robust MDPs
The main results of this section are: (i) the existence of nonrandomized policies minimizing static CVaR (Theorem 3.1), and (ii) the minimal value of static CVaR is the minimal guaranteed expected total discounted payoff for the DM in the RMDP introduced in Chow et al. [4], if the DM plays policies from the MDP; see Corollary 3.4 below. This corollary provides a game-theoretic interpretation of the gap between the optimal value of CVaR and optimal value of this RMDP discovered in Hau et al. [16].
Let us consider an MDP with finite state and action spaces and respectively. The goal is to find a policy minimizing for all and for a given risk level Let
be the optimal value of the static CVaR. The following theorem states the existence of nonrandomized optimal policies minimizing static CVaRs and convergence of finite-horizon optimal values of static CVaRs to the optimal infinite-horizon value. We recall that is fixed. We also recall that throughout the entire paper the discount factor if and if
Theorem 3.1.
For every or for for the static CVaR optimization problem there exists a nonrandomized optimal policy for which for all In addition,
Proof.
We observe that it is sufficient to proof that for every there exists a nonrandomized optimal policy for the initial state Indeed, let be the set of all nonrandomized -horizon policies. If is optimal for an initial state then the nonrandomized policy is optimal, where Let us fix and prove the existence of Let us start with In this case the set is finite. Therefore, the strategic measure for an arbitrary policy can be presented as a convex combination of strategic measures for nonrandomized policies [6, Theorem 1]: there are numbers such that and Since taking CVaR of a mixture of distributions is a concave operation [28, Proposition 3.2],
where such that Thus, for an arbitrary policy we have that which means that that a deterministic policy is optimal for the -horizon problem with the initial state
Let us consider the infinite-horizon problem. In the rest of this proof, let us set for all If we add a constant to the cost function then the all CVaR values for an -horizon problem will be changed by the constant Thus, this addition preserves sets of optimal policies. Since the cost function is bounded, without loss of generality we can assume that the function is nonnegative. Let us assume that and let is an upper bound of the cost function, that is, for all and
Let and Then for and as for all trajectories, where Therefore, for every policy
which implies
and
Let us construct a nonrandomized optimal infinite-horizon policy Let be nonrandomized optimal -horizon policies. Since the sets and where are finite, it is possible to choose a sequence of integers such that for all and for all We define for all
Now let for some there is an increasing sequence of natural numbers such that for each history policies make the same decisions for all We select a subsequence of the sequence such that for each history all policies make the decision at the step which is the same for all By repeating this procedure and taking we construct a nonrandomized policy
Let us show that is indeed an optimal policy. Let us consider a time horizon Then there exists an integer such that the policy is optimal for the horizon and coincides at the first steps with the policy . Therefore, for this policy , (3.1) an be rewritten as
Since the policy is optimal for the horizon
Let Then Since in view of (3.2),
We recall that and denote by the -dimensional Euclidian space. Let us consider the RMDP introduced in Chow et al. [4], which we call the Dynamically augmented RMDP (DRMDP). The DRMDP is defined by a tuple where the state space is action space is uncertainty space is with being the number of states in action sets for the DM at states are uncertainty sets for Nature are
where, and, for
one-step costs for and transition probabilities for and where is the Dirac measure on the interval concentrated at the point A DRMDP is a particular case of the RMDP model described in Appendix A, when the sets set-valued mapping one-step cost function and transition probability have specific forms. In particular, states for DRMDPs are denoted by The sets of policies for the DM and Nature are denoted by and respectively. We shall use notation where and for the expected total discounted payoffs over the horizon with the terminal payoff which is continuous in As explained after formulae (A.2) and (A.4), the optimal values are continuous functions. Since is a finite set, this means that the functions are continuous in
We observe that that is, the set of all policies for the original MDP is the subset of the set of policies for the DM in the DRMDP. We call policies from risk-independent. The fundamental difference between a policy and a policy from is that the policy does not have information about the current and past tail risk levels except the initial tail risk level which we consider to be fixed.
The following theorem states that, for each risk-independent policy for the DM, Nature can play a policy maximizing DM’s losses. The proofs of Theorems 3.2 and A.6 are based on the well-known fact that, if in a two-player stochastic game a policy of one player is fixed, then another player deals with an MDP. In general, the states of this MDP are histories in the game. However, the state space of this MDP can be simplified, if the fixed policy belongs to a special class. For example, if the fixed policy is stationary, then another player deals with an MDP whose states are the states of the stochastic game. There are differences between the proofs of Theorems 3.2 and A.6. The proof of Theorem A.6 follows from the validity of sufficient conditions for the existence of optimal policies for MDPs with setwise continuous transition probabilities [36, 9], which takes place because all the sets are finite. The proof of Theorem 3.2 follows from the validity of sufficient conditions for the existence of optimal policies for MDPs with weakly continuous transition probabilities [36, 8], which takes place because the set and all the sets are finite and because of weak continuity of the transition probability The remaining statements of this section deal with the DRMDP.
Theorem 3.2.
For each risk-independent policy for the DM and for each or there exists a nonrandomized policy which depends on and such that
Proof.
If the DM plays a risk-independent policy Nature deals with an MDP with states The set of available actions at each state is where and The one-step reward for Nature is and this function is bounded and continuous, where If an action is chosen at state then the next state is with the probability where the following conditions hold: (a) (b) (c) (d) Other transitions are impossible since The transition probability is weakly continuous in because is a bounded continuous function and and are distributions on finite sets depending on the finite numbers of conditions. We consider the expected total discounted rewards for this problem with the discount factor The initial state is where which is the initial tail risk level, and This MDP satisfies the weakly continuous conditions, which implies the existence of optimal Markov policies for finite-horizon problems and optimal deterministic policies for infinite-horizon problems [36] or [9, Theorem 2]. These optimal policies define the policy for Nature whose existence is stated in the theorem. ∎
The following theorem establishes the relations between the DRMDP and CVaR of the total discounted reward defined in (2.1). In particular, formula (3.3) illustrates time inconsistency of static CVaR because at some time instance the policy of Nature maximizing the right-hand side of (3.3) may depend on future decisions of the DM playing a possibly nonstationary policy
Theorem 3.3.
For each nonrandomized risk-independent policy for the DM, for each initial tail risk level , and for each or
Proof.
Let us consider the MDP defined at the beginning on the proof of Theorem 3.2 for We fix and compute an optimal policy for Nature. If at time Nature is at state of this MDP, then the optimal 1-step value for Nature is
where there is a nonrandomized optimal 1-step Markov policy [9, Theorem 2], and, in view of Berge’s maximum theorem, the function is continuous. This means that this function is continuous in since all other variables in take a finite number of values. Nature chooses the optimal action without using any knowledge of the previous states and actions. This is true for because the optimal value of does not depend on the distributions of future actions and states following the terminal state
Let us make the induction assumption that for some integer there is a Markov policy such that this policy is optimal for the horizon and the value function is continuous, which means that it is continuous in Then
for where and In view of Berge’s maximum theorem, the function is continuous in and there is a measurable mapping such that the maximum in (3.4) is achieved at [9, Theorem 2] or [10, Theorem 3.3]. Thus, the Markov policy is optimal for the horizon By induction, the Markov policy is optimal for the horizon This Markov policy is defined for Nature in the special MDP. It corresponds to the nonrandomized policy for Nature, which with a small abuse of notations we denote by
If the DM plays the nonrandomized risk-independent policy and Nature plays the nonrandomized policy then, according to Theorem 3.2,
A minor change in the described construction leads to defining an optimal nonrandomized policy for Nature, if the DM plays an arbitrary risk-independent policy but we do not need the explicit form for the policy for Nature if the DM has chosen a possibly randomized risk-independent policy
In view of the CVaR decomposition theorem [29] and the definition of sets for
for
and, for and we have The last equality and (3.5) imply that the theorem is proved for
For
where the first equality follows from unform -a.s. convergence the second equality follows from the proved part of this theorem for and the last one follows from convergence of value iterations for the MDP with weakly continuous transition probabilities [9, Theorem 2] introduced in the proof of Theorem 3.2 when ∎
The following corollary from Theorems 3.1 and 3.3 characterizes the optimal value of static CVaR in terms of the DRMDP. Recall that is the set of policies for the original MDP.
Corollary 3.4.
For every or every , and every
where is the set of nonrandomized risk-independent policies for the DM, and
We recall that the DRMDP satisfies assumptions of an RMDP considered in Appendix A. Thus, the value functions exist and are continuous in for finite and infinite time horizons , and as uniformly in and Theorem A.4 implies the existence of optimal policies and characterizes some classes of optimal policies. In particular, for each player there exist optimal nonrandomized Markov policies, and for each player has a deterministic optimal policy.
Let us consider the gap between the optimal value of CVaR and the value of the RMDP. For or for and for
where the second equality follows from Corollary 3.4 and Theorem A.7, and the inequality follows from It is shown in [16] that this gap can be positive for see also [13].
If then we deal with a one-step problem, which can be solved easily by performing one step of value iterations for the DRMDP with In this case, the gap is equal to 0. Therefore, in the rest of this paper, except Appendix A, we consider
4 Dynamically Augmented CVaR
Let us consider the DRMDP with if and if Theorem 3.3 demonstrates time inconsistency of the static CVaR for MDPs because, for a policy on which CVaR is achieved, current decisions of Nature may depend on future decisions of the DM. The following definition addresses this problem. Optimal decisions of Nature do not depend on future actions of the DM. Therefore, it is natural for Nature to make optimal decisions.
Definition 4.1.
For a policy initial state tail risk level and time horizon or the Dynamically augmented CVaR (DCVaR) is
where is the set of optimal policies for Nature for the -horizon problem.
Since this definitions applies to policies According to Theorem 3.1, for each tail risk level there is an optimal nonrandomized policy minimizing the CVaR for the MDP. If the DM plays a nonrandomized risk-independent policy then, in view of and Theorem 3.3,
Thus, in addition to being a more natural objective function than CVaR, DCVaR can be used for establishing performance guarantees for CVaR. In addition, for the DRMDP
This is true because for each
where the equality follows from the existence and definition of an optimal policy for the DM in the DRMDP, and the inequalities follow from the definition of the DCVaR and from the first equality in Theorem A.7. Thus all inequalities in (4.2) hold in the form of equalities.
According to (A.3), for or and for
We recall that the bounded continuous function is given, the functions are bounded and continuous in and uniformly in as see (A.4). Values can be computed by value iterations, and the interval can be easily discretized for computations. In addition is the unique bounded solution of (4.3) with
Here we would like to mention that time-inconsistency of the static CVaR for MDPs is well-known, and another objective, called the nested CVaR, is broadly used [33, 34, 35, 37, 38, 39, 40]. One of the definitions of the nested CVaR is relevant to abstract dynamic programming [3] dealing with equations for where is an objective function for a finite or infinite time horizon is the state, is the action, and is a terminal cost. For the nested CVaR, the function is CVaR of with respect to a transition probability where is the one-step cost. For an MDP, the optimality equation for the nested CVaR for the risk level is
Expressions (4.3) and (4.4) are closely relevant. In fact, (4.4) can be viewed as (4.3) when the two-dimensional argument of in (4.4) is projected to a single-dimensional argument Also, similar to (4.4), formula (4.3) minimizes CVaR, but the tail risk level depends on the history of the process. As algorithm DCVaR demonstrates, this risk level depends on the value function and previous gains and losses. Contrary to this, for the nested CVaR, the risk level is constant. In this sense, the DCVaR is more flexible than the nested CVaR.
Let us define for or and Then (4.3) becomes
This is the optimality equation for the DRMDP1 defined in the next paragraph. It is easier to deal with the DRMDP1 than with the DRMDP because, as stated in Lemma 7.1, the value functions are concave in when is concave in as assumed in Assumption 5.1.
Let us introduce the DRMDP1 as an RMDP defined by the tuple where the state space the action space uncertainty space sets of available actions and uncertainty sets are the same as in the DRMDP, one-step costs and the transition probability where and Let be the expected total cost for the DRMD1. The following theorem implies that the DRMDP and DRMDP1 correspond to the same problem when where
Theorem 4.2.
For and or
Proof.
The following two corollaries follow directly from Theorem 4.2 and from formula (4.2), which holds in the form of equalities.
Corollary 4.3.
For a finite or infinite horizon initial state and initial risk level , the equality holds for a nonrandomized risk-independent policy for the DM if and only if for every optimal policy for Nature.
Corollary 4.4.
For a finite or infinite horizon let be a nonrandomized risk-independent policy for the DM, and If for every optimal policy for Nature, then
and, in particular, is DM’s optimal policy for the initial state
Formula (4.5) can be rewritten, for and for in the form of (A.1) and (A.2) as
According to (A.5) and (A.6), the sets of optimal actions for the DM are
and for Nature, for
The following theorem provides a sufficient condition under which a policy for the MDP minimizes DCVaR. The assumption that the policy is nonrandomized is not needed. We formulate this assumption only because this theorem is applied to a nonrandomized policies in this paper.
Theorem 4.5.
For a nonrandomized risk-independent policy initial state tail risk level and time horizon or if for every persistently optimal policy for Nature and for every nonnegative integer then formulae (4.6) hold.
Proof.
If then (4.3) for becomes
which is the minimax equation for the sequential deterministic game in which the DM chooses actions, and Nature chooses a transition from the set of transitions having positive probability. The DM is trying to minimize the length of a feasible path while Nature is trying to maximize it. For a horizon a path is a sequence and for it is For a given initial state a path is called feasible if and for all integer values The lengths of finite and infinite-horizon paths are defined by formulae (2.2) and (2.3) for finite and infinite horizon problems respectively. This is a special case of an RMDP considered in Appendix A. In this case, the state space is finite, action sets for the DM are uncertainty sets for Nature are one step costs are and all moves are deterministic to the next states selected by Nature. This problem can be solved easily by value iteration.
We notice that, if then all uncertainty sets can be reduced to singletons where such that if Thus, if then we deal with an MDP with expected total discounted costs.
5 Formulation of the Main Result
Let us consider the DRMDP1. We recall that where the function is continuous in When everywhere in the rest of this paper, except Appendix A, we assume that Assumption 5.1 holds. For example, this assumption holds when the function is nonincreasing in and concave in for all In particular, this assumption holds if the final payoff does not depend on that is,
Assumption 5.1.
The function is concave in
In this section we introduce Algorithm DCVaR constructing for and a nonrandomized policy minimizing among all policies In view of (4.1), this means that The case is addressed in the previous section. Let
We are interesting in dealing with concave functions because, according to Lemma 7.1, the functions and are concave in the tail risk level In addition, if and then the functions and are piecewise linear in
For a real-valued continuous concave function defined on a finite interval we denote by and its right
and left
derivatives respectively, where the concavity of implies If is a real-valued continuous concave function defined on the interval we set and For a function of multiple variables, say with all variables except one, say , being discrete, we shall consider a function in for fixed values of all other parameters. The right and left derivatives in if they exist, will be also denoted as and respectively. If a derivative in exists for some , and it is denoted by which means that We usually apply these notations to functions of two variables where and We recall that, for a real-valued continuous concave function defined on an interval, the right and left derivatives always exist, the function is right-continuous and lower semicontinuous, and the function is left-continuous and upper semicontinuous. A concave function on an interval is differentiable everywhere except for at most a countable set. We also denote by the superdifferential of in at the point In particular, if exists, then We shall also apply this notation for the superdifferential of functions of one and two variables.
For a real number and a real-valued continuous concave function defined on a finite interval in view of the convention and one of the following two possibilities holds:
there is a unique point such that
there is an open interval such that for all
In case (ii) there is a maximal open interval satisfying (ii). If the function is piecewise linear, then in case (i).
Let us consider the sets of optimal actions defined in (4.8). According to Theorem 4.5, if for every optimal policy of Nature, at each finite epoch such that a policy of the DM chooses actions from the state then this policy minimizes in
The following equalities hold in view of Hiriart-Urruty and Lemaréchal [18, p.28]:
In particular, (5.1) and (5.2) imply that,
Algorithm DCVaR for running an optimal policy for an MDP with a finite state space finite action sets available at states transition probability one-step costs and discount factor where if
Inputs: Initial state tail risk level the horizon length or the value functions if or the value function if After an action is selected at the state where is a nonnegative integer such that the next state such that becomes known.
1. Set and choose an arbitrary .
2. Do steps 2.1–2.4 while :
2.1. If then choose an arbitrary number
2.2. Compute:
2.3. If there is a unique point such that then set and Otherwise, set and choose such that is an interior point of the interval, on which the function is linear with the slope in other words, choose such that and there are points and such that .
2.4. If then stop; otherwise, choose an arbitrary optimal action set and then set
If then the algorithm stops after N iterations and returns a finite sentence of actions
implementing an optimal policy at the steps If then the algorithm returns an infinite sequence of actions implementing an optimal policy at the steps and the algorithm can be stopped after a finite number of iterations because the impact of additional steps will be negligibly small due to the discount factor and estimations of stopping times are standard.
As an illustration of Algorithm DCVaR, Figure 1 shows the two main cases addressed by the algorithm. Figure (1(a)) is an example of a case where there is a unique tail risk level being identified in step 2.3, and Figure (1(b)) corresponds to the other case where there is an interval with the desired slope value.
The variable indicates that the tail risk level is either given or identified at step 2.3 as the unique point such that The variable indicates that there is a nonempty open interval on which If then is a piecewise linear function and therefore the existence of the unique described point implies that If and then the functions and are piecewise linear for and Subroutine 1 in Section 6 computes functions
Tail risk levels assigned by Nature are not available to the DM when However, by using formula (5.3), which is based on the analysis of Nature’s optimal policies, at step 2.3 the algorithm detects either the tail risk level or an interval, in which is located, and every optimal action for an internal point of this interval is also optimal if Nature selects another tail risk level from this interval. The main result of this paper is summarized in the following theorem, whose proof is provided in Section 8.
Theorem 5.2.
For or and Algorithm DCVaR generates a nonrandomized risk-independent policy minimizing DCVaR, that is,
6 Properties of the Mass Transfer Problems Solved by Nature
This section studies the problem Nature solves at each horizon or In particular, it studies problem (6.1) motivated by the maximization equation in (4.7). The results of this section are self-contained. Continuity and concavity in of the function in formula (6.1) are assumed because, as shown in the next section, these assumptions are satisfied by the functions and from (4.7). Tneorem 6.2 is the key fact on which formula (5.3) in Algorithm DCVaR is based.
A function where is called piecewise linear, if there is a finite sequence of increasing numbers with and such that the function is affine on each interval
Let be a real-valued function such that for each fixed this function is continuous and concave in For a probability distribution on that is, for all and , let
where .
A function is called feasible for problem (6.1) if for each the function is Borel-measurable on and for all For example, is a feasible function. A feasible function for problem (6.1) is called a solution to problem (6.1) if for all
For each we also consider the set of solutions at In other words, a feasible function is a solution if and only if for all
The following theorem states the existence of solutions and describes the properties of the function
Theorem 6.1.
There exists a solution to problem (6.1), and the function is continuous and concave.
Proof.
Without loss of generality we assume that for all because, if then state can be excluded from the set We observe that is a continuous compact-valued set-valued mapping of into the set of subsets of In addition for all and the function is continuous on Therefore, the Berge maximum theorem implies that are nonempty compact sets for all the function is continuous, and the set-valued mapping is upper-semicontinuous at all In view of the Kuratowsli-Ryll-Nardzewski measurable selection theorem, there is a Borel-measurable optimal solution which implies that all functions are Borel-measurable in Let and where Then Then where the second inequality follows from concavity of ∎
We notice that concavity of implies continuity of but this fact is not used in the proof of Theorem 6.1. The following theorem describes the properties of the solutions to problem (6.1).
Theorem 6.2.
Let be a solution to problem (6.1) and If then for each
In addition, the following relations hold for all
and, if the function is piecewise linear in for each then the function is also piecewise linear.
Proof.
Problem (6.1) can be simplified to a problem which has a natural interpretation. First, without loss of generality we can assume that Indeed, we can always consider the objective function Then all the objective and value functions will be shifted by the constant and the solution sets won’t change. Second, in the rest of this section we assume without loss of generality that for all
Third, let us consider the functions where and vectors where For and let us define Then and problem (6.1) becomes
where
Similarly to problem (6.1), we can consider feasible functions and solutions and solution sets for problem (6.7). Then is a solution (or feasible function) for problem (6.1) if and only if where is a solution (a feasible function, respectively) for problem (6.7).
Thus, in order to prove Theorem 6.1, it is sufficient to prove its statement for problem (6.7) with the functions and replaced with and respectively. In particular, formulae (6.4)–(6.6) become
Problem (6.7) describes the following optimal mass transfer problem. For let us consider intervals in which we call sources. These sources can be interpreted as vertical cylindrical vessels of the same diameter filled with a liquid up to heights Let There is an additional interval which we call the destination. This interval can be interpreted as an empty vertical cylindrical vessel with the same diameter and with height 1. The total value of the liquid at source from level 0 up to the level is where is a concave function with The goal is to fill up the destination by the liquid from sources to maximize for each the total value of the liquid in the destination from the level 0 up to the level . Any amount of the liquid can be taken from any part of each source.
We recall, that for a concave function with and for
and the set is countable. Thus, both functions and where describe the nonlinear unit cost of the liquid at the source depending on the height These unit costs may have arbitrary signs (positive, negative, or 0), but they are nondecreasing functions of the hight since the function is concave in
Let us prove (6.2) and (6.5). If then these formulae hold in the form Let Since the function is concave, the most valuable part of the liquid of the volume at the source is the interval whose value is Thus, if the amounts of liquid should be taken by an optimal policy for the level from sources then the decision to take all the liquid from the interval from each source is optimal, and for all In particular, for So. (6.2) is proved.
Inequality (6.2) can be written as for if In particular, and the strict inequality is impossible. Indeed, if the strict inequality holds, then concavity of and imply the existence of such that for all Since there exists such that Let Then we consider the feasible solution with the same decisions outside of the interval of and allocating all possible resource from the source when the level is in this interval. The formal definition is: for
Let us verify that is a feasible function for problem (6.7). We need to verify the properties of for Let Then In addition, Thus, is feasible, and
which is impossible since is an optimal value, where the last equality in (LABEL:eq6fin) follows from
The proof of (6.6) is similar. In particular, for all is the only solution for and (6.6) holds in the form of So, we consider The same arguments as for (6.2) imply (6.3). Therefore, where is a solution of (6.7) corresponding to a solution of (6.1). The similar argument as in the proof of (6.5) imply that the strict inequality is impossible because otherwise there is a solution such that for some Thus, (6.10) and therefore (6.6) hold. Equations (6.5) and (6.6) imply (6.4). The last claim is correct because, if functions are piecewise linear in for each , then formulae (6.5) and (6.6) imply that the functions and take finite numbers of values when Therefore, the concave function is piecewise linear on the interval ∎
Remark 6.3.
It is possible to construct all solutions of problem (6.1). We do not use these solutions in this paper. So, we do not provide details on constructing them. If then is the unique solution of problem (6.7). If then is the unique solution of (6.7). Thus, is a solution of problem (6.1) for and for this solution is unique. Let Then, let us consider two cases. In case 1, does not belongs to an interval on which the concave function is linear. In case 2, belongs to an interval on which the function is linear. In case 2 we choose in the way that this is the maximal open interval containing on which is linear. In case 1, there is a unique solution at . In case 2, where is the set of all states such that there are points such that and the function is linear on with its derivative on this interval equal to In this case we consider being the maximal open interval satisfying this condition. If then In case 2, and is a solution at iff: (i) for and (ii) and for We notice that, in case 2, if is a singleton, then if and is the unique solution of problem (6.1) at
Remark 6.4.
In order for solutions to have physical meaning of moving the liquid from sources to the destination, the optimal functions should be nondecreasing in To achieve this, it is sufficient to define appropriately the values for when belongs to an interval on which the function is linear, and is not a singleton. In this case, for example, we can move the liquid from sources to the destination sequentially. Sources with smaller numbers can be used first. Such solutions are nondecreasing in Of course, there are other ways to construct nondecreasing solutions.
A concave piecewise linear function on a finite interval can be represented by a finite sequence where are slopes (derivatives at linear intervals) and are lengths of linear intervals where and Formally speaking, if where and
For if a function is represented by a sequence then the function is represented by the sequence with and In view of Theorem 6.2, if all the functions are piecewise linear in then the function is piecewise linear, and it can be constructed in the following way.
Subroutine 1:
1. Merge sequences into a single finite sequence By doing this, merge the intervals with the same slopes. This means that any final set of pairs with distinct indexes and the same slope where and should be replaced with the single pair
3. The resulted finite sequence where and represents the concave piecewise linear function
Subroutine 1 is the major step in recursive computations of value functions for see the following section.
7 Properties of Value Functions and Sets of Optimal Actions.
This section describes the properties of the value functions and for DRMDP1 defined in equations (4.7), where or It also describes the properties of the sets of optimal actions for the DM. We recall that the function is concave and continuous in because the function is continuous in and because Assumption 5.1 holds. We observe that, if then the function is linear in
We observe that for fixed and or the first equation in (4.7) is a particular instance of equation (6.1) with with the variable in (6.1) substituted with with and with
Lemma 7.1.
For and for the functions and defined in (4.7) are continuous and concave in for all and Furthermore, if and is piecewise linear in then these functions are piecewise linear in
Proof.
For we prove this lemma by induction. For , the function is continuous and concave in This is explained in the first paragraph of this section.
Assume that the function is continuous and concave in for some Theorem 6.1 applied to the first formula in (4.7) implies that the functions is continuous and concave in and, in view of the second formula in (4.7), the function is continuous and concave in If the function is piecewise linear in then the last claim in Theorem 6.2 applied to the first formula in (4.7) implies that the functions is piecewise linear in Since for each the set is finite, the second formula in (4.7) implies that the functions are piecewise linear in
In view of (A.4), concave and continuous in functions converge uniformly to Thus, is concave and continuous in ∎
Thus, for each and for each or the function is concave and continuous in on the interval We recall that by definition and . Then for each real number exactly one of the following to possibilities takes place: (i) either there exists a unique such that , or (ii) there exist a unique interval such that on this interval the function is linear with the slope , and if
Lemma 7.2.
Assume that, for or and for some the value function is linear on an interval where Then the following statements hold:
if for some then for all
for
for and
Proof.
According to Lemma 7.1 the functions and are concave in and for all Consider statement (i) and let for some This means that Since and are concave functions, and dominates we have that for all Statements (ii) and (iii) follow from (i). ∎
The state space is infinite, and functions and can be computed by discretizing the interval They can be also computed exactly if the state and action sets and the horizon length are not too large. For each computing can be done by Subroutine 1 described at the end of Section 6. According to the second equation in (4.7), for each functions are lower envelops in of piecewise-linear concave in functions and their computation is a standard task in computational geometry.
8 Proof of Theorem 5.2.
This section contains the proof of Theorem 5.2.
Proof of Theorem 5.2.
For the given time horizon or and for a given sequence Algorithm DCVaR sequentially generates a finite or infinite sequence of optimal actions. Let be the states of the system and be the tail risk levels, and the values are not observed by the DM when In addition to the parameters of the model, the algorithm also uses the value functions , if or it uses the value function if
The parameter takes two values: 1 and 0. means that at the current time epoch the tail risk level is known, and indicates that it is not known. However, in the second case the algorithm finds an interval of possible risk levels, and this is sufficient to choose an optimal action.
At each iteration the algorithm chooses an action . In view of Theorem 4.5, in order to prove the theorem, it is sufficient to show that for all nonnegative integer for every persistently optimal policy played by Nature.
The proof is based on induction in According to step 1, Let an optimal action be selected at an epoch by step 1 or 2.4 of the algorithm. In view of (4.7),
Formula (6.6) implies
where the last inequality holds because and for every Thus,
Similarly to (8.1), formula (6.5) implies
Therefore, in view of inequalities (8.1) and (8.2),
for If there is a unique point such that then is the tail risk level at the state and epoch
If there are multiple points such that that is, the case takes place, then concavity of implies that there is a maximal interval such that the function is linear in and its slope is According to Lemma 7.2, for all the sets coincide, and for Therefore, step 2.1 can be skipped, and future calculations do not depend on the particular value So, in spite of the fact that the DM does not know the states when The algorithm either calculates values or can calculate intervals to which values belong, and optimal action sets coincides for tail risk levels, which are internal points these intervals. Actions from these set are also optimal for the extreme points and of this interval. ∎
9 Extension to Stochastic Cost Functions
In this section, we extend the results of this paper to random one-step cost functions with finite supports. Arbitrary random one-step costs can be approximated by such costs.
In practice, one-step costs can be random. To model random costs, for and let us consider finite sets and real-valued one-step costs where If a control is selected at a state and the system moves to a state a random cost is collected, where has a discrete distribution satisfying for all , and
Let us set Let and for
If we augment the state with the parameter we have a particular case of the original problem with the expanded state space. To explain details, let the augmented state be where is the realization of the random outcome that took place at the previous time instance. At the initial time 0, there are no previous events. In this case, we choose an arbitrary and consider an initial state instead of the original initial state
So, we obtain the MDP with the state space action space sets of available actions transition probabilities
and one-step costs
Notice that the transition probabilities and costs do not depend on Therefore, the value functions also do not depend on the state component In view of this detail, we provide below the minimax equations for DRMDP1 for this model, which are similar to (4.5):
and
where for all and
with where and are the numbers of elements of the sets and respectively.
As a remark, minimax equations (9.1) and (9.2) can be extended to possibly infinite sets and For metric spaces and nonnegative costs equation (9.1) can be rewritten in the integral form
where
and is a set of measurable real-valued functions on Developing specific conditions for correctness of (9.3) for infinite sets and is beyond the scope of this paper.
Acknowledgement
This research was partially supported by the U.S. Office of Naval Research (ONR) under Grant
N000142412608.
The authors thank Jefferson Huang and Pavlo Kasyanov for their comments and express special thanks to Alexander
Shapiro for valuable suggestions and insightful discussions.
Appendix A Results on Robust MDPs Used in this Paper
This appendix provides results on RMDPs used in this paper. In general, for robust optimization problems, some parameters of the model are unknown, and the goal is to select the best solutions for the worst possible values of unknown parameters. Starting from [20] and [27] there are numerous studies of RMDPs motivated by applications. For an RMDP, there is an additional parameter on which one-step costs and transition probabilities depend, and at each step Nature (sometimes called Player II) selects this parameter from a set of available parameters. The DM (sometines called Player I) chooses actions by knowing the past and current states and the past actions chosen by the DM and by Nature. Nature chooses actions by knowing this information and, in addition, by knowing the current action chosen by the DM. Such models were also studied under the names of two-person zero-sum games with perfect information (see e.g., [1, 12, 21]) and minimax control [14]. In RMDP papers [20, 27] the unknown parameter is the transition probability, and one-step costs depend only on states and actions selected by the DM. However, this specification, probably motivated by applications, does not simplify the theory needed in this paper. The models studied in [20, 27] are currently called the models satisfying the property of -rectangularity. Another property called -rectangularity was introduced in [44], and there are other rectangularity properties; see [13, 24, 15, 43]. Models with -rectangularity are relevant to two-person zero-sum stochastic games [21, 41], when the second player does not know the current action selected by the first player.
For the purposes of this paper, it is sufficient to consider an RMDP with a state space being a compact subset of a Euclidean space, with a finite action space for the DM, with action sets of Nature being compact subsets of a Euclidean space, with continuous one-step cost functions, and with weakly continuous transition probabilities. We consider a slightly more general model in this appendix.
An RMDP is defined by a tuple Here is the state space, and there are two players: the DM and Nature. There are two decision sets and where is the set of actions for the DM, and is the set of actions for Nature. We assume that , and are nonempty Borel subsets of Polish (complete, separable, metric) spaces. The sets where are assumed to be nonempty and finite, and are the sets of actions available to the DM at states For each pair where and a set of feasible actions of Nature is defined. We assume that the sets are nonempty and compact subsets of for all and for all We also assume that is a Borel subset of and the set-valued mapping is continuous. Continuity of implies Borel measurability of According to the Arsenin-Kunugui measurable selection theorem, each of the graphs and contains at least one Borel measurable function. This fact, which is needed for the existence of policies, also follows from the Kuratowski and Ryll-Nardzewski measurable selection theorem.
The one-step cost depends on the current state action chosen by the DM, action chosen by Nature, and the next state The function is assumed to be bounded and continuous. The transition probability from to is weakly continuous in Thus, and is a continuous function on for every bounded continuous function
At each time step Nature knows the action chosen by the DM at time when Nature makes a decision. Let be the sets of histories that can be observed by the DM at time The set of histories, that can be observed by Nature at time is A policy of the DM is a sequence of regular transition probabilities from to such that Nature’s policy is a sequence of regular transition probabilities from to such that Let and be the sets of policies for the DM and Nature respectively. Similarly to MDPs, an initial state and a pair of policies define the probability on the space of trajectories An expectation with respect to this probability is denoted by
It is also possible to consider nonrandomized, Markov, and deterministic policies. A policy for the DM is called nonrandomized if for each there exists a mapping such that for all Such a policy is also denoted by A Markov policy for the DM is a sequence of measurable functions such that for all A deterministic policy for the DM is a Markov policy such that for all and all
A policy for Nature is called nonrandomized, if for each there exists a mapping such that for all Such a policy is also denoted by A Markov policy for Nature is a sequence of measurable functions such that for all A deterministic policy for Nature is a Markov policy such that for all and all
If an initial state is and the DM and Nature play policies and respectively, then the expected total finite-horizon payoff of the DM to Nature is
where is a bounded continuous function (the terminal payoff), and for the infinite horizon
For and let us define sequentially
where, in view of Berge’s maximum theorem, each function is continuous, and therefore is a continuous function. Equations (A.1) and (A.2) imply the minimax equation
Let Then all continuous functions are uniformly bounded by where Banach’s fixed point theorem applied to the space of uniformly bounded continuous functions on implies that the minimax operator applied in (A.3) to the function has a unique fixed point which is a bounded continuous function, and as In particular
Let be defined in (A.1) with Thus, (A.2) holds also for and is a unique bounded continuous function satisfying (A.3) or (A.1), (A.2) with
For and for let us consider the following nonempty sets of actions for the DM
minimizing the right-hand side of (A.2), and for Nature, for and
maximizing the right-hand side of (A.1). The sets are finite. As follows from the Berge maximum theorem, the sets are compact.
For an -horizon problem with or a pair of policies is called an equilibrium if for all policies and
Theorem A.1.
For a given horizon or if two pairs of policies and are equilibria, then and are equilibria too, and for all and
Proof.
The definition of an equilibrium implies that for all This equality and the definition of equilibria imply the conclusions of the theorem. ∎
If a pair is an equilibrium for some then is called an optimal policy for the DM, and is called an optimal policy for Nature.
In order to formulate structural properties of optimal policies, we need to introduce additional notations and definitions. Let us fix a horizon or
Let the DM and Nature reset the clock to 0 at time Let the history prior to time be , and let and be the histories the DM and Nature observe respectively during the following units of time after the clock is reset to 0. If the clock is not not reset, then at time the history would be for the DM and for Nature.
Let and the policies for the DM and for Nature, if the clock is reset to 0 at time and the finite sequence of past states and actions is observed. For
are the policies the DM and Nature will use respectively following their original policies and if the clock is reset to 0 at time
Definition A.2.
A policy for the DM ( for Nature) is called perfectly optimal for an -horizon problem, if it is optimal and, for each positive integer and for each the policy () is optimal for the horizon
Definition A.3.
A policy for Nature is called persistently optimal for an -horizon problem, if
for all nonnegative integers all and all
We have that a nonrandomized policy for Nature is persistently optimal if and only if for all nonnegative integers all and all A Markov policy for Nature is persistently optimal if and only if for all nonnegative integers all and all For a deterministic optimal policy for Nature is persistently optimal if and only if for all and all
Theorem A.4.
For every horizon or the following statements hold:
a policy for the DM is perfectly optimal if and only if for all nonnegative integers and for all
a persistently optimal policy for Nature is perfectly optimal;
there exists a Markov perfectly optimal policy for the DM, and a Markov policy for the DM is perfectly optimal if and only if for all nonnegative integers and for all
there exists a Markov persistently optimal policy for Nature, and is a Markov persistently optimal policy for Nature if and only if for all nonnegative integers all and all
if and are optimal policies for the DM and Nature respectively, then
for there exist deterministic optimal policies for the DM and Nature, and all deterministic optimal policy for the DM and for Nature are perfectly optimal;
for a deterministic policy for the DM is optimal if and only if for all and every deterministic persistently optimal policy for Nature is optimal.
Proof.
Let Let be a policy for the DM satisfying the conditions stated in and let be a persistently optimal policy for the Nature. Let us prove that the policies and are perfectly optimal for the DM and Nature respectively. For a history let us denote by its part that does not include the last state, that is, We also define for all policies. Straightforward backward calculations imply that if for and for all then for all In particular, (A.7) holds when If and then backward calculations imply that
Thus, and are perfectly optimal. So, statement and the sufficient condition in are proved. Since (A.7) holds for persistently optimal policies and considered above, Theorem A.1 implies that (A.7) holds for any pair of equilibrium policies. Thus (iii) is proved.
Let us prove the necessary condition in Let be a persistently optimal policy for the DM, and it does not satisfy the conditions stated in Let be a persistently optimal policy for Nature. In view of , be a perfectly optimal policy for Nature. Let us consider the largest nonnegative integer such that for some Then backward calculations imply that and thus the policy is not perfectly optimal. This contradiction implies the necessary condition in Statement follows from when is a Markov policy. Statement follows from definitions of persistently optimal policies and Markov policies for Nature.
Thus, statements are proved for Let us prove them for To prove let us consider a policy for the DM such that for all and all We also choose a persistently optimal policy for Nature for Let us consider a finite-horizon problem with the terminal value Then finite-horizon values for this problem with horizon are equal to and the sets and for this problem are equal to and respectively. Statements for a finite horizon implies that is perfectly optimal for the DM, and is perfectly optimal for Nature for all However, uniformly in , and this implies that policies and are perfectly optimal for the DM and Nature respectively for In particular, statement (iii) holds. To prove the necessary condition in let be a perfectly optimal policy for the DM. Then is perfectly optimal for every finite-horizon problem described in this paragraph. Thus, for all and all In view of Theorem A.1, statements and imply In particular, a deterministic policy for the DM )is perfectly optimal if and only if for all A deterministic policy for Nature is persistently optimal if and only if for all ∎
Corollary A.5.
For or if is an optimal policy for Nature, and is a policy for the DM satisfying the condition for all nonnegative integers then the following statements hold:
(ii) for all nonnegative integers and there exists a persistently optimal policy for Nature such that for all nonnegative integers and
Proof.
(i) Let be a perfectly optimal policy for the DM. Then, in view of Theorem A.4, for all and
Let us define the perfectly optimal policy for the DM: for nonnegative integer and for
Since for all nonnegative integer the probability measures and coincide. Therefore,
(ii) Let be a persistently optimal policy for Nature. Then, in view of Theorem A.4, for all
Let us define the persistently optimal policy for Nature: for nonnegative integer for and for
Let assume that it is possible that for some nonnegative integer Then
Thus, while the DM plays the perfectly optimal policy the optimal policy for Nature can be improved on the set of that has a positive probability if Nature switches to the policy Standard arguments imply that which is impossible since and are optimal policies. The last equality in (ii) follows from (i) since a persistently optimal policy is optimal. ∎
Theorem A.6.
For every horizon or and for every policy for Nature, there is a nonrandomized policy for the DM such that
Proof.
The following theorem presents the minimax equation. We recall that, unlike the case of stochastic games with simultaneous decisions, these equations are not symmetric since Nature knows current decisions of the DM.
Theorem A.7.
For every horizon or
Proof.
Let us fix and let and be optimal policies for the DM and Nature whose existence is stated in Theorem A.4. Let us fix As follows from the definition of optimal policies,
where the inequalities follow from the properties of infimums. Thus, all inequalities in (A.10) are equalities, and the infimum in (A.10) is the minimum since it is achieved at The first equality in (A.9) is proved. In addition,
where the equality follows from the definition of an optimal policy, the first inequality follows from the properties of supremums and from Theorem A.6, the second inequality follows from the inequality between and and the last inequality is taken from (A.10). Thus, these inequalities hold in the form of equalities, the infimum is the minimum, the first supremum is the maximum, and the second equality in (A.9) is proved. ∎
The following statement generalizes Theorem A.7.
Corollary A.8.
Let or and let and be subsets of policies for the DM and Nature such that there exist an optimal policy for the DM and an optimal policy for Nature. Then
Proof.
We observe that, since an optimal policy for
Therefore, similarly to the proof of Theorem A.7, for
and